# Surface Tension of a Topological Phase

### v2

@@ III.2 Role of breaking particle-hole symmetry (class A model)
$H=-\kappa\sum_{i}s_{i}{\bf{\Psi}}_{i}^{\dagger}\sigma_{z}~{}{\bf{\Psi}}_{i}-\sum_{\langle ij\rangle}({\bf{\Psi}}_{i}^{\dagger}{\large{\eta}}_{ij}~{}{\bf{\Psi}}_{j}+{\rm{h.c.}})-\lambda\sum_{i}{\bf{\Psi}}_{i}^{\dagger}(\sigma_{x}+\sigma_{y})~{}{\bf{\Psi}}_{i}-J\sum_{\langle ij\rangle}s_{i}s_{j}-h\sum_{i}s_{i},$ (S23)
where $\lambda$ is the hopping parameter between $A$ and $B$ sites within the same unit cell which breaks the particle-hole symmetry. For uniform $s_{i}$ throughout the system, we assume $M\equiv\kappa s_{i}$ and the Hamiltonian in momentum space for $k_{x},k_{y}\in[-\pi,\pi)$ turns out to be
@@
For $\lambda\neq 0$, the particle-hole symmetry of the Hamiltonian is broken, i.e., $\sigma_{x}H^{*}_{k_{x},k_{y}}\sigma_{x}\neq-H_{-k_{x},-k_{y}}$. When $\lambda\neq 0$, the system with uniform $s_{i}$ hosts topological phases with $C=\pm 1$ for $0<|M|<2\sqrt{1-\lambda^{2}}$ and trivial phase ($C=0$) for $|M|>2\sqrt{1-\lambda^{2}}$ (see Fig. S5(a)). Even in absence of particle-hole symmetry ($\lambda\neq 0$), the ground state energy of the fermionic system with uniform $s_{i}$ remains symmetric under $M\to-M$, i.e., $E_{g}(M)=E_{g}(-M)$. Thus, the sudden quench from $h<0$ to $h>0$ causes no additional correction to bulk energy density ($\Delta\epsilon\to 0$), while there is a substantial correction to surface tension $\Delta\sigma>0$ in topological situation (see Fig. S5(b,c)). Therefore, the enhancement ratio $\gamma=n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}=(1+(\Delta\sigma/\sigma_{\rm{cl}}))^{2}$ for $\lambda\neq 0$ shows similar behavior as for $\lambda=0$ (see Fig. S5(d)) and establishes the increase in critical cluster-size even when particle-hole symmetry is absent.
IV Role of Thermal Fluctuations
@@
In this section we discuss the effects of thermal fluctuations on the surface tension and critical droplet size both for the fermions and for the spin degrees of freedom.
IV.1 Fermionic fluctuations
Figure S6: Surface tension and critical cluster-size at finite temperature: (a) Quantum correction to surface tension $\Delta\sigma$ as a function of $\kappa$ at temperatures $T=0$ and $T=1/\beta=0.33$, (b) Surface tension $\sigma_{\rm{cl}}$ of classical Ising model ($\kappa=0$) as a function of $T$, (c) $\gamma=n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}=(1+(\Delta\sigma/\sigma_{\rm{cl}}))^{2}$ as a function of $\kappa$ at temperatures $T=0$ and $T=1/\beta=0.33$, where $n_{c}^{\rm{cl}}$ is the critical cluster-size for classical Ising model ($\kappa=0$). For all the plots, number of unit cells is $L^{2}=4096$. The parameters chosen are: $J=0.2$ (where $T_{c}\approx 0.45$) and $h=0.02$. Here, classical contribution to surface tension $\sigma_{\rm{cl}}=2J$ at $T=0$. At finite $T$ in (b), $n_{c}^{\rm{cl}}$ is evaluated using Monte Carlo simulation for classical Ising model ($\kappa=0$) via Eq. (S26) and $\sigma_{\rm{cl}}$ is then obtained from $n_{c}^{\rm{cl}}$ using Eq. (S27). In (c), $n_{c}^{\rm{cl}}=145$ for $T=1/\beta=0.33$.
@@ IV.1 Fermionic fluctuations
In order to distill the role of thermal fluctuation on fermions we freeze the spin configurations before and after quench to where all $s_{i}=-1$ and where a region of size $n=l^{2}$ has $s_{i}=1$. However fermions are now assumed to equilibrate to a finite temperature $T$. The correction to surface tension at a finite temperature $\Delta\sigma(T)$ is again computed from $\Delta E(T)=4\Delta\sigma(T)\sqrt{n}$, where the change in energy $\Delta E(T)$ for the fermionic problem at temperature $T=1/\beta$ (where Boltzmann constant is set to unity) is given by
$\Delta E(T)=\frac{\rm{Tr}(e^{-\beta H_{\rm{final}}}~{}H_{\rm{final}})}{\rm{Tr}(e^{-\beta H_{\rm{final}}})}-\frac{\rm{Tr}(e^{-\beta H_{\rm{initial}}}~{}H_{\rm{initial}})}{\rm{Tr}(e^{-\beta H_{\rm{initial}}})}.$ (S25)
@@
Here, $H_{\rm{initial}}$ is the Hamiltonian for fermionic problem before the quench where all $s_{i}=-1$ and $H_{\rm{final}}$ is the Hamiltonian after the quench where there is a square-shaped cluster of size $n=l^{2}$ having $s_{i}=1$ within the region having $s_{i}=-1$. We find that at a finite temperature $T$, the correction to the surface tension $\Delta\sigma(T)$ does not change significantly from its zero temperature ($T=0$) value in both the topological ($20<|\kappa|<$ 2) and trivial ($|\kappa|>2$) situations, as evident from Fig. S6(a). This is expected since the thermal fluctuations at $T<T_{c}$ are perturbative compared to the fermionic bandwidth.
IV.2 Entropic contributions to spin fluctuations
In order to estimate the role of thermal fluctuations on $\sigma_{\rm{cl}}$ which equals $2J$ at $T=0$, we perform the Monte Carlo simulations to evaluate the critical cluster-size $n_{c}^{\rm{cl}}$ at $\kappa=0$ via
$\nu(n_{c}^{\rm{cl}})=\frac{P_{+}(n_{c}^{\rm{cl}})-P_{-}(n_{c}^{\rm{cl}})}{P_{+}(n_{c}^{\rm{cl}})+P_{-}(n_{c}^{\rm{cl}})}=0$ (S26)
and in turn estimate $\sigma_{\rm{cl}}(T)$ using
$\sigma_{\rm{cl}}(T)=2|h|\sqrt{\frac{n_{c}^{\rm{cl}}}{\pi}},$ (S27)
since any correction to $\sigma_{\rm{cl}}(T)$ for $\kappa=0$ is only due to thermal spin fluctuations and has no fermionic contribution. As shown in Fig. S6(b), finite temperature $T$ reduces $\sigma_{\rm{cl}}$ due to entropic contributions coming from the spin fluctuations in the nucleating region. Given the reduction in the effective surface tension, this illustrates that the spin-fluctuations on the boundary of the nucleating region effectively reduces the stiffness of the domain wall. This is consistent with the analytic estimate of the same (see Ref. [27]).
Interestingly, the effect of thermal fluctuations on fermionic and spin together lead to an increase in the relative enhancement $\gamma=n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}=(1+(\Delta\sigma/\sigma_{\rm{cl}}))^{2}$ due to an effective reduction of the $\sigma_{\rm{cl}}$. This enhancement of $\gamma$ at finite $T$ is more dominant in the topological situation ($20<|\kappa|<$ 2) compared to the trivial situation ($|\kappa|>2$), given the fermionic contribution $\Delta\sigma$ to surface tension is significant in the topological regime (see Fig. S6(c)).
V Monte Carlo Simulation Details
In this section, we discuss the details of Monte Carlo simulation employed to explore the time evolution of the coupled spin-fermion system.
@@ V Monte Carlo Simulation Details
Figure S7: Effective Monte Carlo simulation: (a,b) Local magnetization $s_{x,y}$ for unit cell at position $(x,y)$ where initial cluster-size with up-spins is $n=441$ ($n>n_{c}^{\rm{Q}}$) at times $t=0,20,40$ (in units of Monte Carlo steps per site) using (a) coupled spin-fermion Monte Carlo simulation (Eq. (S29) and Eq. (S30)) and (b) effective Monte Carlo simulation (Eq. (S31) and Eq. (S32)). The parameters chosen are: $\kappa=1.2$, $J=0.2$, $T=1/\beta=0.33$ (where $T_{c}\approx 0.45$), $h=0.02$. (c,d) Change in ground state energy $\Delta E_{i}^{\rm{Q}}$ (see Eq. (S29)) and its comparison with $\Delta E_{i}^{\rm{Q,eff}}$ (see Eq. (S32)) for fermionic system when $i$-th spin is flipped from (c) $s_{i}=1$ to $s_{i}=-1$, (d) $s_{i}=-1$ to $s_{i}=1$. Here, $N_{i}=\sum_{j\in{\mathcal{N}}_{i}}s_{j}$ where $j$ runs over all the four nearest neighbors (${\mathcal{N}}_{i}$) of $i$-th spin. We find that $\Delta E_{i}^{\rm{Q}}\approx\Delta E_{i}^{\rm{Q,eff}}$. Here, we have chosen $\kappa=1.2$ and number of unit cells in square lattice $L^{2}=4096$.
V.1 Coupled spin-fermion Monte Carlo dynamics
In Monte Carlo (MC) simulation for the classical Ising model ($\kappa=0$), a random spin $s_{i}$ is chosen and the energy $\Delta E_{i}$ required to flip the spin (i.e., $s_{i}\to-s_{i}$) is calculated as:
$\Delta E_{i}=2Js_{i}\sum_{j\in{\mathcal{N}}_{i}}s_{j}+2hs_{i},$ (S28)
where $j$ runs over all the nearest neighbors (${\mathcal{N}}_{i}$) of $i$-th spin. Now, if $\Delta E_{i}<0$ or if $p_{i}\leq\exp(-\beta\Delta E_{i})$ for a randomly chosen $p_{i}\in[0,1]$ where $\beta=1/T$, the spin flip is accepted and the $i$-th spin becomes $-s_{i}$. If none of these conditions is satisfied, the spin-flip is rejected and $i$-th spin remains $s_{i}$.
When classical Ising field is coupled to fermionic field ($\kappa\neq 0$), we solve the fermionic problem to compute $\Delta E_{i}$ at each step of Monte Carlo simulation. In this method, we propose a spin flip $s_{i}\to-s_{i}$ at a randomly chosen $i$-th unit cell at each Monte Carlo step. We then explicitly calculate the difference $\Delta E_{i}^{\rm{Q}}$ in the ground state energies between the configurations after and before the proposed $i$-th spin-flip for the fermionic problem as:
$\Delta E_{i}^{\rm{Q}}=E^{\rm{Q}}\left(s_{1},...,-s_{i},...,s_{L^{2}}\right)-E^{\rm{Q}}\left(s_{1},...,s_{i},...,s_{L^{2}}\right),$ (S29)
$\Delta E_{i}=2Js_{i}\sum_{j\in{\mathcal{N}}_{i}}s_{j}+2hs_{i}+\Delta E_{i}^{\rm{Q}}.$ (S30)
Eq. (S30) can be used to check the acceptance or rejection of the spin-flip at each Monte Carlo step when $\kappa\neq 0$. Under this protocol, the spin fluctuations and their effect on fermions are exactly accounted for. An exact MC simulation for $\kappa=1.2$ and initial droplet size $n=441$ is shown in Fig. S7(a), illustrating the growth of the droplet (as $n>n_{c}^{\rm{Q}}$). This method, however, is numerically expensive, since at every MC step, a matrix diagonalization for the fermionic system needs to be performed. In the next section, we show how this step can be simplified by devising an effective Monte Carlo method.
V.2 Effective Monte Carlo simulation and validity
@@
In order to simplify the exact MC method as discussed in the previous section, we here show that we can also calculate $\Delta E_{i}$ approximately for a spin-flip ($s_{i}\to-s_{i}$) when $\kappa\neq 0$ from the quantum correction to surface tension $\Delta\sigma$ as:
$\Delta E_{i}=2Js_{i}\sum_{j\in{\mathcal{N}}_{i}}s_{j}+2hs_{i}+\Delta E^{\rm{Q,eff}}_{i},$ (S31)
@@
$\Delta E^{\rm{Q,eff}}_{i}=(\Delta\sigma)s_{i}\sum_{j\in{\mathcal{N}}_{i}}s_{j},$ (S32)
@@
where $\Delta\sigma$ is evaluated at $T=0$, as shown in the inset of Fig. 2(a) of the main text. This effectively replaces the fermion diagonalization step with an effective correction from the fermions. In Fig. S7(b), we show the effective MC dynamics with the same initial random seed and same initial configurations as shown previously using exact MC. We next estimate the validity of this step by comparing $\Delta E_{i}^{\rm{Q}}$ from each Monte Carlo dynamics (see Eq. (S29)) with effective $\Delta E_{i}^{\rm{Q,eff}}$. The energy correction can be estimated for the two processes $s_{i}=1\rightarrow-1$ and for $s_{i}=-1\rightarrow 1$ for different configurations of neighborhood spins characterized by $N_{i}=\sum_{j\in{\mathcal{N}}_{i}}s_{j}$ that can take values $-4,-2,0,2,4$. In Fig. S7(c,d), we show the variation of $\Delta E^{\rm{Q,eff}}_{i}$ and $\Delta E_{i}^{\rm{Q}}$ for both the processes for different values of $N_{i}$ at $\kappa=1.2$. The values are essentially indistinguishable, establishing the validity and utility of the effective MC procedure. Thus, while exact droplet growth can have minor fluctuations between exact MC dynamics and effective MC dynamics, the effective MC method is particularly apt in evaluating the coarse-grained description of the system.
V.3 Typical and averaged droplet growth
Figure S8: Local magnetization: Local magnetization $s_{x,y}$ for unit cell at position $(x,y)$ where initial cluster-size with up-spins are $n=441$ ($n>n_{c}^{\rm{Q}}$), $n=225$ ($n_{c}^{\rm{Q}}>n>n_{c}^{\rm{cl}}$), $n=81$ ($n_{c}^{\rm{cl}}>n$) at times $t=0$, $t=20$, $t=100$ (in units of Monte Carlo steps per site) when $\kappa=0$ and $\kappa=1.2$. The parameters chosen are: $J=0.2$, $T=1/\beta=0.33$ (where $T_{c}\approx 0.45$), $h=0.02$. Number of unit cells in square lattice is $L^{2}=4096$.
Using the Monte Carlo simulation scheme described above, we compute local magnetization $s_{x,y}$ at location $(x,y)$ and time $t$ (in units of Monte Carlo steps per site or after $tL^{2}$ Monte Carlo steps where $L^{2}$ is the number of spins in square lattice) in a typical realization of the simulation. $s_{x,y}$ for various times $t$ in a typical realization for classical ($\kappa=0$) and topological ($20<|\kappa|<$ 2) situations are depicted in Fig. S8 where the boundary of the droplets can be clearly seen during their growth (shrinkage). As expected, any typical droplet boundary follows a very rugged trajectory.
To estimate the universal behavior of such growth, configuration averaging is performed which has been reported in the main text. An average local magnetization $\langle s_{x,y}\rangle$ where averaging is performed over $10^{3}$ realizations of the simulations is provided in Fig. 3(b) of the main text. Averaging over various realizations in $\langle s_{x,y}\rangle$ renders the growing (shrinking) droplets circularly symmetric.
VI Role of Slow Quench
Figure S9: Dynamics of average local magnetization under slow quench: (a) Linear quench protocol of $h(t)$ as a function of $t$ (see Eq. (S33)), (b,c) Average local magnetization $\langle s_{x,y}\rangle$ for the unit cell with coordinate $(x,y)$ for a cluster of up-spins having initial size $n=441$ at time $t=100$ (in the units of Monte Carlo steps per site) for different values of $\tau$ when (b) $\kappa=0$ and (c) $\kappa=1.2$. In all the plots, the parameters chosen are: $J=0.2$, $T=1/\beta=0.33$ (where $T_{c}\approx 0.45$), $h_{f}=0.02$. The number of unit cells in square lattice is $L^{2}=4096$ and number of simulations considered is $1000$. Cluster size at fixed $t=100$ reduces with increasing $\tau$.

### v3

@@ B.2 Role of breaking particle-hole symmetry (class A model)
$H$ $=-\kappa\sum_{i}s_{i}{\bf{\Psi}}_{i}^{\dagger}\sigma_{z}~{\bf{\Psi}}_{i}-\sum_{\langle ij\rangle}({\bf{\Psi}}_{i}^{\dagger}{\large{\eta}}_{ij}~{\bf{\Psi}}_{j}+{\rm{h.c.}})$
$-\lambda\sum_{i}{\bf{\Psi}}_{i}^{\dagger}(\sigma_{x}+\sigma_{y})~{\bf{\Psi}}_{i}-J\sum_{\langle ij\rangle}s_{i}s_{j}-h\sum_{i}s_{i},$ (36)
where $\lambda$ is the hopping parameter between $A$ and $B$ sites within the same unit cell which breaks the particle-hole symmetry. For uniform $s_{i}$ throughout the system, we assume $M\equiv\kappa s_{i}$ and the Hamiltonian in momentum space for $k_{x},k_{y}\in(-\pi,\pi]$ turns out to be
@@
For $\lambda\neq 0$, the particle-hole symmetry of the Hamiltonian is broken, i.e., $\sigma_{x}H^{*}_{k_{x},k_{y}}\sigma_{x}\neq-H_{-k_{x},-k_{y}}$. When $\lambda\neq 0$, the system with uniform $s_{i}$ hosts topological phases with $C=\pm 1$ for $0<|M|<2\sqrt{1-\lambda^{2}}$ and trivial phase ($C=0$) for $|M|>2\sqrt{1-\lambda^{2}}$ (see Fig. 11(a)). Even in absence of particle-hole symmetry ($\lambda\neq 0$), the ground state energy of the fermionic system with uniform $s_{i}$ remains symmetric under $M\to-M$, i.e., $E_{g}(M)=E_{g}(-M)$. Thus, the sudden quench from $h<0$ to $h>0$ causes no additional correction to bulk energy density ($\Delta\epsilon\to 0$), while there is a substantial correction to the interfacial line energy $\Delta\sigma>0$ in topological situation (see Fig. 11(b,c)). Therefore, the enhancement ratio $\gamma=n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}=(1+(\Delta\sigma/\sigma_{\rm{cl}}))^{2}$ for $\lambda\neq 0$ shows similar behavior as for $\lambda=0$ (see Fig. 11(d)) and establishes the increase in critical cluster-size even when particle-hole symmetry is absent.
Appendix C Role of thermal fluctuations
@@
Now we discuss the effects of thermal fluctuations on the interfacial line energy and critical droplet size both for the fermions and for the spin degrees of freedom.
C.1 Fermionic fluctuations
@@
In order to distill the role of thermal fluctuation on fermions we freeze the spin configurations before and after quench to where all $s_{i}=-1$ and where a region of size $n=l^{2}$ has $s_{i}=1$. However fermions are now assumed to equilibrate to a finite temperature $T$. It is useful to note we work with $T<T_{c}$ where classical transition temperature $T_{c}\approx 2.269J$ [46] (where Boltzmann constant is set to unity).
The correction to the interfacial line energy at a finite temperature $\Delta\sigma(T)$ is again computed from $\Delta E(T)=4\Delta\sigma(T)\sqrt{n}$, where the change in energy $\Delta E(T)$ for the fermionic problem at temperature $T=1/\beta$ (where Boltzmann constant is set to unity) is given by
$\Delta E(T)=\frac{{\rm{Tr}}(e^{-\beta H_{\rm{final}}}~H_{\rm{final}})}{{\rm{Tr}}(e^{-\beta H_{\rm{final}}})}-\frac{{\rm{Tr}}(e^{-\beta H_{\rm{initial}}}~H_{\rm{initial}})}{{\rm{Tr}}(e^{-\beta H_{\rm{initial}}})}.$ (38)
@@
Here, $H_{\rm{initial}}$ is the Hamiltonian for fermionic problem before the quench where all $s_{i}=-1$ and $H_{\rm{final}}$ is the Hamiltonian after the quench where there is a square-shaped cluster of size $n=l^{2}$ having $s_{i}=1$ within the region having $s_{i}=-1$. We find that at a finite temperature $T$, $\Delta\sigma(T)$ does not change significantly from its zero temperature ($T=0$) value in both the topological ($0<|\kappa|<2$) and trivial ($|\kappa|>2$) situations, as evident from Fig. 12(a). This is expected since the thermal fluctuations at $T<T_{c}$ are perturbative compared to the fermionic bandwidth.
@@
Figure 13: Effective Monte Carlo simulation: (a,b) Local magnetization $s_{x,y}$ for unit cell with position $(x,y)$ at times $t=0,20,40$ (in units of Monte Carlo steps per site) using (a) coupled spin-fermion Monte Carlo simulation (Eq. (10) and Eq. (11)) and (b) effective Monte Carlo simulation (Eq. (41) and Eq. (42)). Here, the initial cluster-size with up-spins is $n=441$ ($n>n_{c}^{\rm{Q}}$). The parameters chosen are: $\kappa=1.2$, $J=0.2$, $T=1/\beta=0.33$ (where $T_{c}\approx 0.45$), $h=0.02$. (c,d) Change in ground state energy $\Delta E_{i}^{\rm{Q}}$ (see Eq. (10)) and its comparison with $\Delta E_{i}^{\rm{Q,eff}}$ (see Eq. (42)) for fermionic system when $i$-th spin is flipped from (c) $s_{i}=1$ to $s_{i}=-1$, (d) $s_{i}=-1$ to $s_{i}=1$. Here, $N_{i}=\sum_{j\in{\mathcal{N}}_{i}}s_{j}$ where $j$ runs over all the four nearest neighbors (${\mathcal{N}}_{i}$) of $i$-th spin. We find that $\Delta E_{i}^{\rm{Q}}\approx\Delta E_{i}^{\rm{Q,eff}}$. Here, we have chosen $\kappa=1.2$ and number of unit cells in square lattice $L^{2}=4096$.
C.2 Entropic contributions to spin fluctuations
In order to estimate the role of thermal fluctuations on classical $\sigma_{\rm{cl}}$ which is $\sim 2J$ at $T=0$, we perform the Monte Carlo (MC) simulations. Here, a finite-sized cluster is subjected to the non-equilibrium situation (after a sudden quench of $h$) and the critical cluster-size $n_{c}^{\rm{cl}}$ at $\kappa=0$ is evaluated via
$\nu(n_{c}^{\rm{cl}})=\frac{P_{+}(n_{c}^{\rm{cl}})-P_{-}(n_{c}^{\rm{cl}})}{P_{+}(n_{c}^{\rm{cl}})+P_{-}(n_{c}^{\rm{cl}})}=0$ (39)
where $P_{+}(n)$ and $P_{-}(n)$ are the number of MC steps in which size $n$ of a cluster increases to $(n+1)$ and decreases to (n1)(n-1) respectively [40]. Here, for the cluster-size $n<n_{c}^{\rm{cl}}$, the tendency of shrinkage of the cluster is greater than that of growth, implying $P_{+}<P_{-}$ and $\nu<0$, while for $n>n_{c}^{\rm{cl}}$, the tendency of growth of the cluster is greater than that of shrinkage, implying $P_{+}>P_{-}$ and $\nu>0$. Therefore, $\nu=0$ for $n=n_{c}^{\rm{cl}}$.
From the critical cluster size $n_{c}^{\rm{cl}}$, we then estimate $\sigma_{\rm{cl}}(T)$ using
$\sigma_{\rm{cl}}(T)=2|h|\sqrt{\frac{n_{c}^{\rm{cl}}}{\pi}},$ (40)
since any correction to $\sigma_{\rm{cl}}(T)$ for $\kappa=0$ is only due to thermal spin fluctuations and has no fermionic contribution. As shown in Fig. 12(b), finite temperature $T$ reduces $\sigma_{\rm{cl}}$ due to entropic contributions coming from the spin fluctuations in the nucleating region [17]. Given the reduction in the effective $\sigma_{\rm{cl}}$ with $T$, this illustrates that the thermal fluctuations on the boundary of the nucleating region effectively reduces the stiffness of the domain wall.
Interestingly, the effect of thermal fluctuations on fermionic and spin together lead to an increase in the relative enhancement $\gamma=n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}=(1+(\Delta\sigma/\sigma_{\rm{cl}}))^{2}$ due to an effective reduction of the $\sigma_{\rm{cl}}$. This enhancement of $\gamma$ at finite $T$ is more dominant in the topological situation ($0<|\kappa|<2$) compared to the trivial situation ($|\kappa|>2$), given the fermionic contribution $\Delta\sigma$ is significant in the topological regime (see Fig. 12(c)).
Figure 14: Dynamics of average local magnetization under slow quench: (a) Linear quench protocol of $h(t)$ as a function of $t$ (see Eq. (43)), (b,c) Average local magnetization $\langle s_{x,y}\rangle$ for the unit cell with coordinate $(x,y)$ for a cluster of up-spins having initial size $n=441$ at time $t=100$ (in the units of Monte Carlo steps per site) for different values of $\tau$ when (b) $\kappa=0$ and (c) $\kappa=1.2$. In all the plots, the parameters chosen are: $J=0.2$, $T=1/\beta=0.33$ (where $T_{c}\approx 0.45$), $h_{f}=0.02$. The number of unit cells in square lattice is $L^{2}=4096$ and number of simulations considered is 10001000. Cluster size at fixed $t=100$ reduces with increasing $\tau$.
Appendix D Effective Monte Carlo simulation and validity
@@
In order to simplify the exact MC method, we here show that we can also calculate $\Delta E_{i}$ approximately for a spin-flip ($s_{i}\to-s_{i}$) when $\kappa\neq 0$ from the quantum correction to the interfacial line energy $\Delta\sigma$ as:
$\Delta E_{i}=2Js_{i}\sum_{j\in{\mathcal{N}}_{i}}s_{j}+2hs_{i}+\Delta E^{\rm{Q,eff}}_{i},$ (41)
@@
$\Delta E^{\rm{Q,eff}}_{i}=(\Delta\sigma)s_{i}\sum_{j\in{\mathcal{N}}_{i}}s_{j},$ (42)
@@
where $\Delta\sigma$ is evaluated at $T=0$. This effectively replaces the fermion diagonalization step with an effective correction from the fermions. In Fig. 13(a,b), we show the exact and the effective MC dynamics with the same initial random seed and same initial configurations for $\kappa=1.2$. We next estimate the validity of this step by comparing $\Delta E_{i}^{\rm{Q}}$ from each Monte Carlo dynamics (see Eq. (10)) with effective $\Delta E_{i}^{\rm{Q,eff}}$. The energy correction can be estimated for the two processes $s_{i}=1\rightarrow-1$ and for $s_{i}=-1\rightarrow 1$ for different configurations of neighborhood spins characterized by $N_{i}=\sum_{j\in{\mathcal{N}}_{i}}s_{j}$ that can take values 4,2,0,2,4-4,-2,0,2,4. In Fig. 13(c,d), we show the variation of $\Delta E^{\rm{Q,eff}}_{i}$ and $\Delta E_{i}^{\rm{Q}}$ for both the processes for different values of $N_{i}$ at $\kappa=1.2$. The values are essentially indistinguishable, establishing the validity and utility of the effective MC procedure. Thus, while exact droplet growth can have minor fluctuations between exact MC dynamics and effective MC dynamics, the effective MC method is particularly apt in evaluating the coarse-grained description of the system.
Appendix E Role of Slow Quench